Find the perimeter of a rectangle if three of its vertices are (5, −2), (−3, −2), and (−3, 3).
Answer by chessace(471)

The side along x axis is (5 - (-3)) = 8.
'' y axis (3 - (-2)) = 5.
perimeter(retangle) = 2 * sum of sides = 2 * 13 = 26